# mO Revanth[#1]

### v1

@@ Abstract
The growing collaboration between humans and AI models in generative tasks has introduced new challenges in distinguishing between human-written, AI-generated, and human-AI collaborative texts. In this work, we collect a multilingual, multi-domain, multi-generator dataset FAIDSet. We further introduce a fine-grained detection framework FAID to classify text into these three categories, meanwhile identifying the underlying AI model family. Unlike existing binary classifiers, FAID is built to capture both authorship and model-specific characteristics. Our method combines multi-level contrastive learning with multi-task auxiliary classification to learn subtle stylistic cues. By modeling AI families as distinct stylistic entities, FAID offers improved interpretability. We incorporate an adaptation to address distributional shifts without retraining for unseen data. Experimental results demonstrate that FAID outperforms several baseline approaches, particularly enhancing the generalization accuracy on unseen domains and new AI models. It provide a potential solution for improving transparency and accountability in AI-assisted writing.
@@ 1 Introduction
AI has progressively evolved from an assistant tool to a creator or initiator, from helping polish papers and generate code to initiate proposals and draft essays, while humans increasingly serve as optimizers and reviewers. In such deeply collaborative human-AI settings, measuring human contribution becomes challenging, yet clarifying authorship is critical for accountability and transparency, particularly in educational and academic contexts Wang et al. (2025). This work aims to identify human involvement by distinguishing the origin of a given text: fully AI-generated, fully human-written, or collaboratively produced, in multilingual context.
@@
Numerous studies have explored multilingual AI-generated text detection, but most focus either on binary detection: human vs. AI Wang et al. (2024d, b), or on fine-grained detection limited to English Abassy et al. (2024); Koike et al. (2024); Wang et al. (2023); Zhang et al. (2024a). Moreover, both struggle with generalization to unseen domains, languages, and new AI generators Wang et al. (2024c); Li et al. (2024).
@@
Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain and multi-generator dataset for fine-grained detection, i.e., identifying a text into three categories: AI-generated, human-written, and human-AI collaborative; and (ii) introducing a detection framework FAID to ameliorate poor generalization performance. [^1] Our code and data are accessible at github.com/ngocminhta/FAID.
@@
Our dataset focuses on academic field including domains of research paper abstracts, student theses, and scholarly reports, and involve recent LLM families, e.g., GPT, Gemini, DeepSeek and Llama OpenAI (2024); Team et al. (2024); DeepSeek-AI (2024); Dubey et al. (2024).
@@
FAID treats each AI model as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features. This was achieved by optimizing a language encoder with multi-level author relationship (e.g., stylistic similarity between texts from the same LLM family is greater than the similarity between a human and an AI) to capture author-specific distinguishable signals of an input text based on contrastive learning, along with the task of fine-tuning a classifier to recognize its origin. This multitask learning process forces the encoder to reorganize the hidden space, representations of texts from the same authors distributed closer than from different authors.
@@
Practically, given that generations produced by models developed by the same company tend to have similar writing styles due to similar model architecture, training data and training strategy (more details in Appendix C), we consider each AI family as an author. This can also help the detector acquire prior knowledge of future LLMs from the same company. Experiments show that FAID consistently outperforms other baseline detectors in both in-domain and out-of-domain evaluations. Our contributions can be summarized as follows:
@@
  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained AI-generated text detection with 84k examples.
@@
  • We propose a detection framework FAID to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in hidden representation.
@@
  • Experimental results show that FAID outperforms baseline detectors, particularly over unseen domains and generators. Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to others with ground truth authorship in the database, implicitly enhancing interpretability with cues explaining why a text is assigned to a specific class.
2 Backgrounds
Advanced LLMs have changed how we create and refine text. Instead of originating every word, humans now more often act as post-editors or reviewers, stepping in after AI drafts the initial content. This collaborative workflow spans fields from academic publishing and journalism to education and social media, making hybrid human-AI texts the new norm Cheng et al. (2024); Lee et al. (2022).
@@ 2 Backgrounds
Such a shift calls question into binary authorship detection: human vs. AI. Texts can be fully human-written, fully AI-generated, or collaborative text (e.g., human-written, AI-polished; human-initiated, AI-continued; AI-generated, human-edited; and deeply mixed text) Artemova et al. (2025). This demands the fine-granularity authorship detection, and even tracing back to a specific model family, to measure human contributions Hutson (2025). Then, clear attribution of AI involvement helps uphold honesty in research and preserve scientific integrity. Responding to these challenges, we collect a new dataset, and propose a detection framework which generalizes well to new domains, languages, and models, showing high accuracy and reliability.
@@ 2.1 Fine-grained AI-generated Text Datasets
Many prior studies have explored fine-grained AI-generated text datasets across various forms of human-AI collaboration as shown in Table 3, for example, MixSet (Zhang et al., 2024b), and Beemo (Artemova et al., 2025). More datasets and detailed information regarding the label space, and the coverage of domains, languages and LLMs for each are summarized in Table 3.
Despite these contributions, they have been limited to English. There is a noticeable gap in the availability of large-scale, fine-grained multilingual AI-generated text detection datasets. To bridge this gap, we collect a multilingual fine-grained AI-generated text detection dataset. It encompasses 84k texts generated by the latest LLMs and includes diverse forms of human-AI collaborative generations. This dataset can facilitate the development of more robust and generalizable detection models that are capable of handling complex multilingual collaborative scenarios.
@@ 2.2 Generalization of AI-generated Text Detection
Recent studies such as M4GT Wang et al. (2024c) and LLM-DetectAIve Abassy et al. (2024) have highlighted a persistent challenge for both binary and fine-grained AI-generated text detection: poor generalization to unseen domains, languages, and generators. Many detection methods show a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors to deal with real-world scenarios and evolving LLMs outputs.
Various approaches have been proposed to improve generalization. OUTFOX Koike et al. (2024) leverages adversarial in-context learning, dynamically generating challenging examples to enhance robustness, but still faces limitations in domain transferability and computational efficiency. LLM-DetectAIve Abassy et al. (2024) adopts fine-grained multi-way classification and incorporates domain-adversarial training to reduce overfitting, yet its generalization to truly unseen domains and generators remains limited. Also, its current version lacks multilingual support. SeqXGPT Wang et al. (2023) focused on sentence-level detection by combining log-probability features with convolutional and self-attention mechanisms, which helps capture subtle mixed-content signals and improves generalization across input styles. However, its reliance on specific model features and limited semantic representation constrains adaptability to new generators and domains.
A recent representative work, the DeTeCtive framework Guo et al. (2024), introduces multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, though it primarily focuses on binary classification and does not fully address human-AI collaborative texts.
@@ 2.3 Contrastive Learning for AI-Generated Text Detection
Contrastive learning has been extensively used to improve the sentence representation by pulling semantically similar sentences closer and pushing dissimilar ones apart, reorganizing the hidden space of sentence representation by semantics. For example, a representative work SimCSE regarded an input sentence under dropout noise as its semantically similar counterpart (i.e., positive pairs), and trained the encoder to close their distance. Additionally, they natural language inference (NLI) pairs, treating entailment pairs as positives and contradiction pairs as hard negatives (Gao et al., 2021). For negative pairs with contradictory meanings, the embeddings were trained to be distant in the latent space. Similarly, DeCLUTR Giorgi et al. (2021) constructed positive pairs by extracting different spans from the same texts, and negative pairs are sampled from different texts.
We adopt this philosophy to reorganize the latent space based on writing styles, pulling human-written texts to cluster together while remaining distant from AI-generated ones. Analogous to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement in collaborative texts, ranging from 0 (fully AI-generated) to 1 (fully human-written).
DeTeCtive also took advantage of contrastive learning and was used in binary AI-generated text detection Guo et al. (2024). Based on a multi-task framework, DeTeCtive was trained to learn writing-style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs. AI) to capture distinguishable signals. For inference, the pipeline encoded the input text as a hidden vector and applied dense retrieval to match the cluster, depending on the stylistic similarity with previously indexed training feature database. Additionally, instead of retraining the model when faced with new data, they encoded new data with the trained encoder to obtain embeddings, and then added them to the feature database to augment. This largely improved the generalizability in unseen domains and new generators.
However, this approach only considers two types of text: human-written and AI-generated, without further consideration of human-AI collaborative texts. Our work fills this gap with the goal of improving generalization performance for fine-grained AI-generated text detection.
@@ 3 FAIDSet
To fill the gap of fewer publicly-available multilingual fine-grained AI-generated text datasets, we collected a multilingual, multi-domain and multi-generator AI-generated text dataset FAIDSet. It contains texts generated by LLMs, written by humans and collaborated by both humans and LLMs, resulting in a total of $\sim$84k examples. See Appendix B for more details about FAIDSet.
FAIDSet covers two domains: student theses and paper abstracts, where identifying authorship is critical, across two languages including Vietnamese and English. We collected student theses from the database of Hanoi University of Science and Technology, and paper abstracts from arXiv and Vietnam Journals Online [^2] https://vjol.info.vn/ (VJOL).
Models and Label Space. We used the following multilingual LLMs families to produce both AI and human-AI collaborative texts: GPT-4/4o, Llama-3.x, Gemini 2.x, and Deepseek V3/R1. Regarding the human-AI collaborative text, we include AI-polished, AI-continued, and AI-paraphrased where the models are requested to polish or paraphrase inputs while ensuring the accuracy of any figures and statistics.
Diverse Prompt Strategies. We generate data with diverse tones and contexts while ensuring the accuracy of the content. Depending on the data source and context, we craft prompts to create varied outputs suitable for different real-world application scenarios. We generate responses with different tones using prompts such as "You are an IT student…" and "…who are very familiar with abstract writing…", which are described more in Appendix B.2.
Quality Control. Given that the task of AI-generated text detection naturally has gold labels when collecting data, we perform quality control by randomly sampling 10-20 instances for each domain, source, and generative model in data collection, and manually assess whether or not the generated texts are fluent and logically reasonable, to avoid incoherence, disorganization, or illogical texts. Such cases introduce remarkably distinguishable signals from human-written text, affecting the data quality. We adjust either the prompts or the generation parameters to address these problems.
@@ 4 Methodology
In this section, we first clarify task definition and then elaborate our detection method.
4.1 Task Definition
In this work, we formulate the task as a three-class classification problem: human-written, AI-generated, and human-AI collaborative texts. The human-AI collaborative category involves a range of interaction between humans and AI systems, such as human-written, AI-polished; human-initiated, AI-continued; human-written, AI-paraphrased; AI-generated, AI-humanized, etc.
Given the growing variety and complexity (e.g., deeply mixed text) of collaborative patterns, we do not exhaust all patterns. Instead, we consolidate all forms of human-AI collaboration into a single label to maintain practical simplicity and model generalizability. This approach reflects real-world usage, where using AI tools to enhance clarity or expression is increasingly common and often ethically acceptable.
We also investigate the dataset and find that the AI models within the same family tend to have similar writing styles and text distributions, due to their shared training data and architecture (see Appendix C). Therefore, we consider each model’s family to be an author with unique writing styles.
@@ 4.2 Framework Overview
Figure 1: Training architecture. Leveraging multi-level contrastive learning loss, we fine-tune a language model (we select XLM-RoBERTa Vamvas and Sennrich (2023), see Appendix D) based on the human, human-AI and AI texts, to force the model to reorganize the hidden space, pulling the embeddings within the same author families closer, and pushing the embeddings from different authors farther. We train an encoder that can represent text with distinguishable signals to discern authorship of text.
We aim to enable the encoder to capture multi-level similarities between authors (AI, human and human-AI) as Eq (6), forcing model to learn features that can differentiate distinct authors in a high-dimensional hidden vector space. We denote $\mathrm{S}_{c}$ as cosine similarity, $\phi(\cdot)$ is the encoder function and $P_{i}$, $P_{j}$, ($51\leq i\leq j<$ 5) are distributions of different text subsets. We have:
@@
where $P_{1}$ corresponds to the distribution generated by a particular LLM family, $P_{2}$ to the distribution generated by any LLMs, $P_{3}$ to the distribution of a collaborative text generated by humans and the LLM family of $P_{1}$, $P_{4}$ to the distribution of collaborative text generated by humans and any LLM families, and $P_{5}$ to the distribution of human-written text.
To clarify the rationale for configuring FAID to expect that the similarity of a text $x$ (from lower-level distributions $P_{1}$ or $P_{2}$) with samples from $P_{3}$ is generally greater than or equal to its similarity with samples from $P_{4}$, consider the following.
  • If $x\in P_{1}$ ($x$ is generated by a particular LLM): Let $y_{\textit{LHS}}$ be drawn from $P_{3}$ and $y_{\textit{RHS}}$ be drawn from $P_{4}$. Naturally, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ should be greater than $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. This is because $P_{3}$ contains texts that share a direct LLM family origin with $x$.
  • If $x\in P_{2}$ ($x$ is generated by any LLM): In this scenario, with $y_{\textit{LHS}}$ from $P_{3}$ and $y_{\textit{RHS}}$ from $P_{4}$ as defined above, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is generally expected to be equal to $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. Since $x$ can originate from any LLM, it does not inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
This configuration aims to ensure that closeness in distribution corresponds to heightened similarity after encoding, encouraging the model to discern fine-grained multi-level relations.

### v2

# FAID: Fine-Grained AI-Generated Text Detection Using Multi-Task Auxiliary and Multi-Level Contrastive Learning

@@ Abstract
The growing collaboration between humans and AI models in generative tasks has introduced new challenges in distinguishing between human-written, LLM-generated, and human–LLM collaborative texts. In this work, we collect a multilingual, multi-domain, multi-generator dataset FAIDSet. We further introduce a fine-grained detection framework FAID to classify text into these three categories, and also to identify the underlying LLM family of the generator. Unlike existing binary classifiers, FAID is built to capture both authorship and model-specific characteristics. Our method combines multi-level contrastive learning with multi-task auxiliary classification to learn subtle stylistic cues. By modeling LLM families as distinct stylistic entities, we incorporate an adaptation to address distributional shifts without retraining for unseen data. Our experimental results demonstrate that FAID outperforms several baselines, particularly enhancing the generalization accuracy on unseen domains and new LLMs, thus offering a potential solution for improving transparency and accountability in AI-assisted writing.
@@ 1 Introduction
LLMs have evolved from an assistant tool to a creator or initiator, from helping polish papers and generate code to initiate proposals and draft essays, while humans increasingly serve as optimizers and reviewers. In such deeply collaborative human–LLM settings, measuring human contribution becomes challenging, yet clarifying authorship is critical for accountability and transparency, particularly in educational and academic contexts Wang et al. (2025). This work aims to identify human involvement by distinguishing the origin of a given text (in a multilingual context): (a) fully LLM-generated, (b) fully human-written, or (c) collaboratively produced.
@@
Numerous studies have explored multilingual LLM-generated text detection, but most have focused either on binary detection, i.e., human vs. LLM Wang et al. (2024d, b), or on fine-grained detection limited to English Abassy et al. (2024); Koike et al. (2024); Wang et al. (2023); Zhang et al. (2024a). Moreover, both struggle with generalization to unseen domains, languages, and LLM generators Wang et al. (2024c); Li et al. (2024).
@@
Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain and multi-generator dataset, FAIDSet, for fine-grained detection, i.e., identifying a text into three categories: LLM-generated, human-written, and human–LLM collaborative; and (ii) introducing a detection framework FAID to improve generalization performance.
@@
Our dataset FAIDSet focuses on the academic field including paper abstracts, student theses, and scholarly reports, and contains generation by a variety of families of LLMs, e.g., GPT, Gemini, DeepSeek, and Llama OpenAI (2024); Team et al. (2024); DeepSeek-AI (2024); Dubey et al. (2024).
@@
Our detection framework FAID treats each LLM as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features. We achieve this by optimizing a language encoder with multi-level author relationship (e.g., the stylistic similarity between texts from the same LLM family is greater than the that between a human and an LLM) to capture author-specific distinguishable signals of an input text using contrastive learning, along with the task of fine-tuning a classifier to recognize the input text’s origin. This multitask learning process forces the encoder to reorganize the hidden space, so that the representations of texts by the same authors are distributed closer than such by different authors.
@@
In practice, given that generations produced by LLMs from same company tend to have similar writing styles due to similar model architecture, training data, and training strategy (see Appendix D for more detail), we consider each LLM family as an “author.” This can also help the detector acquire prior knowledge of future LLMs from the same company. Our experiments show that FAID consistently outperforms other baseline detectors in both in-domain and out-of-domain evaluations. Our contributions can be summarized as follows:
@@
  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained LLM-generated text detection with 84k examples.
@@
  • We propose a detection framework FAID to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in the hidden representation.
@@
  • We show that FAID outperforms baseline detectors, particularly over unseen domains and generators. Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to other texts from our database, labeled with ground truth authorship labels.
2 Background and Related Work
Advancements in LLMs have fundamentally reshaped the process of text production and refinement. Rather than originating every word independently, people increasingly assume the roles of post-editors or reviewers, intervening after an initial draft has been generated by an LLM. This collaborative paradigm extends across diverse domains, including academic publishing, journalism, education, and social media, thus making hybrid human–AI authorship an emerging norm Cheng et al. (2024); Lee et al. (2022).
@@ 2 Background and Related Work
Such a shift calls question into binary authorship detection: human vs. AI. Texts can be fully human-written, fully AI-generated, or collaborative text (e.g., human-written, AI-polished; AI-generated, human-edited; and deeply mixedArtemova et al. (2025). Addressing these concerns necessitate fine-granularity authorship detection, and even tracing back to a specific LLM family, in order to assess the extent of human contribution Hutson (2025). Ensuring transparent disclosure of LLM involvement is critical for upholding research honesty and preserving the integrity of research. In response to these challenges, we collect a new dataset, FAIDSet, and propose a detection framework, FAID, which generalizes well to new domains, languages, and models, achieving consistently high accuracy and reliability.
@@ 2.1 Fine-Grained AI-generated Text Datasets
Many prior studies have explored fine-grained AI-generated text datasets across various forms of human–AI collaboration, e.g., MixSet (Zhang et al., 2024b), and Beemo (Artemova et al., 2025). See Table 3 in the Appendix, for a discussion of more datasets and detailed information regarding the label space, and the coverage of domains, languages and LLMs for each one.
However, all these studies were limited to English. A substantial gap remains in the availability of large-scale, fine-grained multilingual LLM-generated text detection datasets. To bridge this gap, we collect a multilingual fine-grained LLM-generated text detection dataset, which encompasses 84k texts generated by the latest LLMs and includes diverse forms of human–LLM collaborative generations. This dataset can facilitate the development of more robust and generalizable detection models that are capable of handling complex multilingual collaborative scenarios.
@@ 2.2 Generalization of AI-Generated Text Detection
A recent study, M4GT-Bench, Wang et al. (2024c) has highlighted a persistent challenge for both binary and fine-grained AI-generated text detection: poor generalization to unseen domains, languages, and generators. Many detection methods have shown a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors to deal with real-world scenarios and evolving LLMs outputs.
Various approaches have been proposed to improve generalization. OUTFOX Koike et al. (2024) leverages adversarial in-context learning, dynamically generating challenging examples to enhance robustness, but still faces limitations in domain transferability and computational efficiency. LLM-DetectAIve Abassy et al. (2024) adopts fine-grained classification and incorporates domain-adversarial training to reduce overfitting; yet, its generalization to unseen domains and generators remains limited; moreover, its current version lacks multilingual support. SeqXGPT Wang et al. (2023) focuses on sentence-level detection by combining log-probabilities with convolutional and self-attention mechanisms, which helps to capture subtle mixed-content signals and improves generalization across input styles. However, its reliance on specific model features and limited semantic representation constrains adaptability to new generators and domains.
Finally, the DeTeCtive framework Guo et al. (2024), introduces multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, but it primarily focuses on binary classification and does not fully address human–AI collaborative texts.
@@ 2.3 Contrastive Learning for AI-Generated Text Detection
Contrastive learning has been extensively used to improve sentence representation by pulling semantically similar sentences closer and pushing dissimilar ones apart. SimCSE treats a sentence subjected to dropout noise as a semantically similar counterpart (i.e., a positive pair) and trains the encoder to minimize the distance between the original and the noise sentence. It further leverages natural language inference pairs, considering entailment pairs as positives and contradiction pairs as hard negatives (Gao et al., 2021). It is then trained to maximize the separation between negative pairs in the embedding space. DeCLUTR Giorgi et al. (2021) constructed positive pairs by extracting different spans from the same texts, and by sampling negative pairs from different texts.
We adopt the same core idea, but reorganize the latent space based on writing style, pulling human-written texts to cluster together while remaining distant from LLM-generated ones. Similarly to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement in collaborative texts, ranging from 0 (fully LLM-generated) to 1 (fully human-written).
Another work, DeTeCtive Guo et al. (2024), also took advantage of contrastive learning and was used in binary AI-generated text detection. Based on a multi-task framework, it was trained to learn writing-style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs. AI) to capture distinguishable signals. For inference, their pipeline encoded the input text as a hidden vector and applied dense retrieval to match the cluster, depending on the stylistic similarity with respect to a database of previously indexed training features. Additionally, instead of retraining the model when faced with new data, they encoded new data with the trained encoder to obtain embeddings, and then added them to the feature database to augment. This largely improved the generalizability to unseen domains and new generators.
However, this approach distinguishes only between two categories of text (human-written vs. LLM-generated) while overlooking the increasingly prevalent class of human–LLM collaborative texts. Our work bridges this gap with the aim of enhancing generalization performance for fine-grained LLM-generated text detection.
@@ 3 FAIDSet
We collected a new multilingual, multi-domain and multi-generator LLM-generated text dataset, FAIDSet. It contains texts generated by LLMs, written by humans and collaborated by both humans and LLMs, resulting in a total of 84k examples; see Appendix B for more details about FAIDSet.
FAIDSet covers two domains: student theses and paper abstracts, where identifying authorship is critical, across two languages (Vietnamese and English). We collected student theses from the database of Hanoi University of Science and Technology, and paper abstracts from arXiv and Vietnam Journals Online [^1] https://vjol.info.vn/ (VJOL).
Models and Label Space. We used the following multilingual LLM families to produce LLM and human–LLM collaborative texts: GPT-4/4o, Llama-3.x, Gemini 2.x, and Deepseek V3/R1. Regarding the human–LLM collaborative text, we include LLM-polished, LLM-continued, and LLM-paraphrased, where the models are requested to polish or paraphrase inputs while ensuring the accuracy of any figures and statistics.
Diverse Prompt Strategies. We generated data with diverse tones and contexts while ensuring the accuracy of the content. Depending on the data source and context, we crafted prompts to create varied outputs suitable for different real-world scenarios. We generated responses with different tones using prompts such as “You are an IT student…” and “…who are very familiar with abstract writing…”. See Appendix B.2 for the full list. While FAIDSet does not capture the full natural diversity of in-the-wild LLMs outputs, the controlled setup enables us to systematically model stylistic and collaborative signals across multiple languages and families. This makes FAIDSet both a reproducible training resource and a benchmarking corpus: it provides reliable supervision for developing detectors while also serving as a testbed for assessing generalization to more diverse, unseen scenarios.
Quality Control. To avoid bad machine-generated texts, which can introduce remarkably distinguishable signals, we performed quality control by randomly sampling 10-20 instances for each domain, source, and LLM generator. Manual inspection focused on fluency, coherence and factual plausibility. Overall, the generated texts demonstrated high linguistic quality with most outputs being fluent and logically reasonable. Nevertheless, occasional issues such as repetitive phrasing, incomplete reasoning or overuse of formal expressions were observed. In that cases, we adjusted the prompts (e.g., by specifying desired length or style) or refined generation parameters to improve diversity and logical consistency. After these adjustments, the quality across generators and domains was found to be stable and satisfactory.
@@ 4 Methodology
Task Definition
Our task is a three-class classification problem: human-written vs. LLM-generated vs. human–LLM collaborative text detection.
The human–LLM collaborative category involves a range of interaction between humans and LLM systems, such as human-written, LLM-polished; human-initiated, LLM-continued; human-written, LLM-paraphrased, etc. Given the growing variety and complexity (e.g., deeply mixed text) of collaborative patterns, this is not exhaustive. Instead, we consolidate all forms of human–LLM collaboration into a single label to maintain practical simplicity and model generalizability. This reflects real-world usage, where using LLM tools to enhance clarity or expression is increasingly common and often ethically acceptable.
Our analysis of the dataset revealed that the LLM models within the same family tend to have similar writing style and text distributions, due to their shared training data and architecture (see Appendix D). Thus, we consider each model’s family to be an “author” with a unique writing style.
@@ 4.1 Framework Overview
Figure 1: Training architecture. Leveraging multi-level contrastive learning loss, we fine-tune a language model (we select XLM-RoBERTa Vamvas and Sennrich (2023), see Appendix E) based on the human, human–LLM and LLM-generated texts, to force the model to reorganize the hidden space, pulling the embeddings within the same author families closer, and pushing the embeddings from different authors farther. We train an encoder that can represent text with distinguishable signals to discern authorship of text.
Here, we shift the focus to the detector, an encoder-based model that forms the core of FAID. The detector encodes each text into a high-dimensional embedding space to quantify cross-source similarity, enabling the study of both intra- and inter-family relationships. It is trained to capture multi-level similarities between authors by learning a representation space where closely related sources (e.g., LLMs from the same family) form tighter clusters, while dissimilar ones (e.g., human vs. LLM) are pushed farther apart.
Let $\mathrm{S}_{c}$ be cosine similarity, $\phi(\cdot)$ be the encoder function, and $P_{i}$, $P_{j}$, ($1\leq i\leq j<5$) be the distributions of different text sources. We aim for the model to encode representations that satisfy the following constraint:
@@
where $P_{1}$ corresponds to the distribution generated by a particular LLM family, $P_{2}$ is the distribution generated by any LLM, $P_{3}$ is the distribution of collaborative text generated by human and a LLM family of $P_{1}$, $P_{4}$ is the distribution of collaborative text generated by humans and any LLM families, and $P_{5}$ is the distribution of human-written text.
To clarify the rationale for configuring FAID to expect that the similarity of a text $x$ (from lower-level distributions $P_{1}$ or $P_{2}$) with samples from $P_{3}$ is generally greater than or equal to its similarity with samples from $P_{4}$, consider the following:
  • If $x\in P_{1}$ ($x$ is generated by a particular LLM): Let $y_{\textit{LHS}}$ be drawn from $P_{3}$ and $y_{\textit{RHS}}$ be drawn from $P_{4}$. Naturally, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is greater than $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. This is because $P_{3}$ contains texts that share a direct LLM family origin with $x$.
  • If $x\in P_{2}$ ($x$ is generated by any LLM): Here, with $y_{\textit{LHS}}$ from $P_{3}$ and $y_{\textit{RHS}}$ from $P_{4}$ as defined above, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is generally expected to be equal to $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. Since $x$ can originate from any LLM, it doesn’t inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
This configuration aims to ensure that closeness in distribution corresponds to higher similarity after encoding, thus encouraging the model to discern fine-grained multi-level relations.